# Localising the Seizure Onset Zone from Single-Pulse Electrical Stimulation Responses with a Transformer

### v1

@@ Abstract
Epilepsy is one of the most common neurological disorders, and many patients require surgical intervention when medication fails to control seizures. For effective surgical outcomes, precise localisation of the epileptogenic focus – often approximated through the Seizure Onset Zone (SOZ) – is critical yet remains a challenge. Active probing through electrical stimulation is already standard clinical practice for identifying epileptogenic areas. This paper advances the application of deep learning for SOZ localisation using Single Pulse Electrical Stimulation (SPES) responses. We achieve this by introducing Transformer models that incorporate cross-channel attention. We evaluate these models on held-out patient test sets to assess their generalisability to unseen patients and electrode placements.
Our study makes three key contributions: Firstly, we implement an existing deep learning model to compare two SPES analysis paradigms—namely, divergent and convergent. These paradigms evaluate outward and inward effective connections, respectively. Our findings reveal a notable improvement in moving from a divergent (AUROC: 0.574) to a convergent approach (AUROC: 0.666), marking the first application of the latter in this context. Secondly, we demonstrate the efficacy of the Transformer models in handling heterogeneous electrode placements, increasing the AUROC to 0.730. Lastly, by incorporating inter-trial variability, we further refine the Transformer models, with an AUROC of 0.745, yielding more consistent predictions across patients. These advancements provide a deeper insight into SOZ localisation and represent a significant step in modelling subject-specific intracranial EEG electrode placements in SPES. Future work will explore integrating these models into clinical decision-making processes to bridge the gap between deep learning research and practical healthcare applications.
Data and Code Availability
For this study, we utilised an open-source dataset (van Blooijs et al., 2023a), which comprises electrocorticography (ECoG) recordings from 74 patients aged between 4 and 51 years undergoing evaluation for epilepsy surgery. We restricted our analyses to the 35 patients for whom each electrode was labelled as within or outside the SOZ. The dataset is available at https://openneuro.org/datasets/ds004080. Further details can be found in Section 4.1. Data preprocessing and model training code are available at https://github.com/norrisjamie23/Localising_SOZ_from_SPES.
Institutional Review Board (IRB)
As the dataset used in this study is open-source, this study does not require IRB approval.
@@ 1 Introduction
Epileptic seizures originate from diverse brain networks, exhibiting highly individualised, time-varying interictal and ictal patterns that often require uniquely personalised treatments in order to achieve complete seizure control. Around 1 in 100 people will be diagnosed with epilepsy (Vaughan et al., 2018), and approximately one-third of cases are resistant to medication (Kalilani et al., 2018). Many of these patients are candidates for surgery, which is a safe and effective treatment option to reduce or altogether eliminate seizures (Wiebe et al., 2001). This approach entails identifying and removing focal brain regions implicated in seizure generation, with the seizure onset zone (SOZ) being the primary clinically identifiable proxy for these critical areas. However, identifying the SOZ requires intracranial recordings of spontaneously occurring seizures. Seizures are comparatively rare events and often require extended hospital stays to capture, incurring high costs and stress for patients. Even with the current gold standard evaluation, outcomes remain comparatively poor in cases requiring invasive intracranial evaluation. Moreover, a putative SOZ is not identified at all in approximately 15% of cases, preventing surgical intervention (UK Children’s Epilepsy Surgery Collaboration et al., 2021).
@@
SPES elicits two primary types of responses: early and delayed. Early responses, emerging within 100ms post-stimulation and time-locked to the stimulus, are termed cortico-cortical evoked potentials (CCEPs) (Figure LABEL:fig:paradigmsD). These CCEPs signify directed or effective connectivity; for instance, if site A is stimulated, a CCEP observed at site B indicates directed neural propagation from A to B. While CCEPs occur in epileptogenic and non-epileptogenic tissue, their characteristics reflect the proximity of the stimulation and recording sites to the SOZ (Hays et al., 2023). However, the common practice of averaging CCEPs across repeated trials (typically 10-50) to enhance the signal-to-noise ratio risks obscuring significant inter-trial variability. This variability could be a crucial indicator of epileptogenicity in specific brain regions (Cornblath et al., 2023).
In contrast, delayed responses, which occur between 100ms and 1s post-stimulation and resemble epileptiform discharges, reflect focal cortical irritability or even epileptogenicity (Figure LABEL:fig:paradigmsB). These responses have been reported in patients with temporal and frontal lobe epilepsies, and the removal of sites exhibiting these is associated with overall favourable outcomes (Valentín et al., 2005). Unlike the consistent and time-locked nature of CCEPs, delayed responses are neither consistent across trials nor time-locked to the stimulus, meaning they are less apparent when averaging across trials.
@@
Recent advances in machine learning, particularly in deep learning, are well-positioned to model complex multidimensional time-series data, such as those generated by SPES. This capability makes deep learning methods particularly promising for integrating multi-channel SPES responses to localise the SOZ. However, determining the most effective approach to leverage these advanced techniques remains a significant challenge due to particular data features, such as subject-specific electrode placement and variability in the background EEG activity. This leads to several crucial questions for refining and enhancing these methodologies needing to be answered.
@@
A second area of exploration involves integrating information across multiple SPES trials. Classifying individual trials results in numerous predictions for each site. Opting to average across trials addresses this issue. Whilst this averaging process enhances the signal-to-noise ratio, it risks obscuring delayed responses and other important forms of cross-trial variability. Consequently, there is a need for a method that integrates information across trials and effectively factors in this variability.
Lastly, effectively modelling subject-specific and heterogeneous multi-channel data presents a significant challenge. Most machine learning applications to time-series data assume a fixed number of channels with consistent ordering. However, intracranial EEG data varies in channel numbers and placements across patients, impeding model generalisability. Therefore, developing methods that can accommodate the unique and complex nature of subject-specific intracranial EEG datasets is an essential issue that needs to be addressed.
This paper introduces Transformer models with cross-channel attention aimed at enhancing generalisability to unseen patients and managing the complexities of intracranial EEG data. Our study focuses on:
@@
  • Implementing and evaluating Transformer models, outperforming existing methods with an AUROC of 0.730.
  • Refining the Transformer models to incorporate inter-trial variability, further increasing the AUROC to 0.745, demonstrating enhanced robustness and precision.
The rest of this paper is organised as follows: We first review related work on SOZ localisation from SPES and Transformers in intracranial EEG. This is followed by our methodology, detailing the various models employed in this study. The experiments section then introduces the dataset and outlines the comparative analysis. This is followed by the results section, which presents our findings. The paper concludes with a discussion of these findings and our concluding remarks.
@@ 2.1 Localising the SOZ from SPES responses
Malone et al. (2022) evaluated various machine learning algorithms for localising the SOZ from SPES responses, classifying single-channel responses based on whether the recording electrode was in the SOZ. The most effective methods included a 1D convolutional neural network, a gradient-boosted decision tree, and an ensemble approach. In a different study, Yang et al. (2024) focused on computing various CCEP features to classify both stimulated and recording electrodes as SOZ/non-SOZ. These methods, however, consider each channel in isolation, leading to multiple predictions per channel and thereby not integrating the multidimensional data features into the classification. In contrast, Johnson et al. (2022) adopted a multi-channel approach using a divergent paradigm with a ResNet convolutional neural network. This method classifies the stimulated stereo-EEG (SEEG) channel based on responses from other channels. A limitation of this approach is that the network architecture chosen for this analysis requires a consistent number of channels and associates parameters with specific channel positions. To mitigate this, the authors randomly selected a subset of 40 channels for each pass, sorting these channels based on their proximity to the stimulated electrode. This method’s generalisability was validated using leave-one-out cross-validation. While this approach is suited to extracting temporal features, its effectiveness could benefit from adopting a more sophisticated method to focus on important response channels. Rather than having parameters associated with each of a fixed number of channels, capturing a context-aware representation across all channels may deal better with the heterogeneity of subject-specific seizure networks. Furthermore, this fixed channel count necessitates balancing the model’s prediction accuracy and generalisability.
@@ 3.1 Notation
To describe each network, we use the following notation. The output for a given stimulation trial is denoted as $x_{d}\in\mathbb{R}^{N\times C_{d,i}\times T}$ for the divergent paradigm, and $x_{c}\in\mathbb{R}^{N\times C_{c,i}\times T}$ for the convergent paradigm. Here, $N$ represents the number of trials, $C_{d,i}$ (or $C_{c,i}$ for a convergent paradigm) indicates the number of channels for patient $i$, and $T$ is the number of time points. Averaging these responses across trials results in $\bar{x}\in\mathbb{R}^{C_{i}\times T}$, which enhances the signal-to-noise ratio and facilitates a single prediction per electrode. Similarly, the standard deviation across trials is represented as $S_{x}\in\mathbb{R}^{C_{i}\times T}$, capturing the variability across trials.
@@ 3.2 Models
    Transformerbase
    An encoder-only Transformer (Vaswani et al., 2017) with cross-channel attention. A convergent paradigm was used.
  4. 4.
    Transformerall
    An extension to Transformerbase with additional input features.
@@
Figure 2: Network architectures. A: CNN. A ResNet creates a cross-channel embedding, and a multi-layer perceptron (MLP) classifies this as being in or out of the Seizure Onset Zone (SOZ). B: Encoder-only Transformer with cross-channel attention. C: Embedding for Transformerbase. D: Embedding for Transformerall.
@@ 3.2.3 Transformerbase
These embeddings are then subjected to spatial attention, facilitating integration across channels. This process learns a representation across all channels, drawing on the context provided by their interactions. This strategy overcomes the limitations of fixed channel inputs and enables the model to weigh channels based on their features without making assumptions based on the channel order.
3.2.4 Transformerall
To overcome certain limitations of Transformerbase, we developed an extended version, Transformerall (Figure LABEL:fig:modelsD). We made several modifications to enhance SOZ localisation capabilities:
Firstly, given the potential relationship between inter-trial variability and epileptogenicity (Cornblath et al., 2023) and that delayed responses are inconsistent across trials, we hypothesised that averaging across trials might obscure crucial data. To address this, we incorporated standard deviation across trials ($S_{x_{c}}$) alongside the mean ($\bar{x_{c}}$) as part of the model’s input. Each channel was represented as a 2D input to the CNN, combining both time series.
Secondly, since CNNs have some degree of shift-invariance, time-locked responses like CCEPs might not be optimally captured. To mitigate this, we introduced an additional branch in our per-channel embedding. Specifically, the first 300ms of both $\bar{x_{c}}$ and $S_{x_{c}}$ – corresponding to the CCEP component – were directly concatenated to the CNN output (Figure LABEL:fig:modelsD). By skipping the CNN, the linear layer directly embeds the CCEP, which has no shift-invariance.
Lastly, the physical distance between stimulating and response sites was identified as a potentially informative feature for SOZ localisation. For instance, large CCEP amplitudes observed over varying distances, not just between nearby channels, could indicate epileptogenicity. Therefore, we included the distance between recording and response sites as an additional feature in the model.
@@ 4.1 Dataset
Details of the dataset are included in the Data and Code Availability section. The ECoG data were recorded at a sampling rate of 2048 Hz, and these patients exhibited a wide range of seizure onset locations, spanning various regions across the brain. Stimulation was performed at 0.2 Hz with a current intensity of 4mA or 8mA. For additional information on the dataset, including the number and distribution of electrode locations, see Appendix A.
@@ 4.2 Dataset split
We employed repeated k-fold cross-validation with $k=5$ and 5 repeats. To assess model generalisability to new patients and ECoG electrode locations, each patient was present in one fold, ensuring no patient data overlap between the training and testing sets. Further details are provided in Appendix B.
@@ 4.3 Preprocessing
Data were first bandpass filtered to 1-150 Hz in line with previous work in the same task (Johnson et al., 2022). Further preprocessing was applied in line with previous work using the same dataset (van Blooijs et al., 2023b), as outlined in Appendix C.
@@ 4.4 Evaluation criteria
To assess model performance across repeated k-fold cross-validation, we employed both threshold-independent and threshold-dependent metrics, each computed per patient and then averaged across patients, folds, and repeats. We used the Area Under the Receiver Operating Characteristic (AUROC) and the Area Under the Precision-Recall Curve (AUPRC) for threshold-independent metrics. We report the specificity, sensitivity, and Youden’s index (Youden, 1950) for threshold-dependent metrics, which consider binary rather than continuous predictions. We also report the p-value testing for an increase in model performance from CNNconvergent to Transformerbase. Further details are included in Appendix D.
@@ 4.5 Hyperparameter tuning
We undertook a hyperparameter search for each model using Optuna (Akiba et al., 2019) to ensure that observed differences in model performance were not due to inadequate hyperparameter configurations. This process led to the selection of 49 channels for CNNdivergent and 37 channels for CNNconvergent. Further details of the hyperparameter search spaces and the chosen hyperparameters for each model are in Appendix E.
@@ 4.6 Channel sensitivity analysis
Although the rationale for adopting a Transformer over a CNN is its potential to extract features from relevant channels more effectively, it is worth considering the extent to which the performance of Transformerbase could be attributed to the availability of more channels during inference. If this were the case, it would indicate that cross-channel attention is not the cause of these improvements. To investigate this aspect, we undertook a sensitivity analysis with Transformerbase, focusing on how its performance depends on the number of channels available during inference.
Table 1: Model Performance Metrics. Mean and standard deviation of the AUROC, AUPRC, specificity, sensitivity, and Youden’s index for each model across all test folds and runs. The baseline AUPRC, based on the prevalence of the positive class, is 0.136.
Model AUROC AUPRC Specificity Sensitivity Youden
CNNdivergent $0.574\pm 0.065$ $0.204\pm 0.056$ $0.598\pm 0.205$ $0.500\pm 0.199$ $0.098\pm 0.105$
CNNconvergent $0.666\pm 0.067$ $0.305\pm 0.074$ $0.671\pm 0.147$ $0.532\pm 0.214$ $0.202\pm 0.121$
Transformerbase $0.730\pm 0.082$ $0.377\pm 0.119$ $0.636\pm 0.185$ $0.597\pm 0.267$ $0.233\pm 0.147$
Transformerall $0.745\pm 0.098$ $0.378\pm 0.121$ $0.642\pm 0.181$ $0.663\pm 0.212$ $0.305\pm 0.143$
4.7 Transformer performance in Frontal/Temporal SOZs
One of the motivations for including the standard deviation across trials in Transformerall was the potential to better identify delayed responses, which the averaging process could obscure. Since these are observed in the frontal and temporal lobes (Valentín et al., 2005), the patient cohort was divided into two groups for this analysis: those primarily with frontal or temporal lobe SOZs and those without. We computed the AUROC scores for both groups for Transformerbase and Transformerall. The underlying hypothesis was that if the observed improvement in model performance were primarily due to enhanced detection of delayed responses, the improvement would be particularly pronounced in the subcohort of patients with a higher incidence of frontal and temporal lobe SOZs.

### v2

# Localising the Seizure Onset Zone from Single-Pulse Electrical Stimulation Responses with a CNN Transformer

@@ Abstract
Epilepsy is one of the most common neurological disorders, often requiring surgical intervention when medication fails to control seizures. For effective surgical outcomes, precise localisation of the epileptogenic focus – often approximated through the Seizure Onset Zone (SOZ) – is critical yet remains a challenge. Active probing through electrical stimulation is already standard clinical practice for identifying epileptogenic areas. Our study advances the application of deep learning for SOZ localisation using Single-Pulse Electrical Stimulation (SPES) responses, with two key contributions. Firstly, we implement an existing deep learning model to compare two SPES analysis paradigms: divergent and convergent. These paradigms evaluate outward and inward effective connections, respectively. We assess the generalisability of these models to unseen patients and electrode placements using held-out test sets. Our findings reveal a notable improvement in moving from a divergent (AUROC: 0.574) to a convergent approach (AUROC: 0.666), marking the first application of the latter in this context. Secondly, we demonstrate the efficacy of CNN Transformers with cross-channel attention in handling heterogeneous electrode placements, increasing the AUROC to 0.730111https://github.com/norrisjamie23/Localising_SOZ_from_SPES. These findings represent a significant step in modelling patient-specific intracranial EEG electrode placements in SPES. Future work will explore integrating these models into clinical decision-making processes to bridge the gap between deep learning research and practical healthcare applications.
@@ 1 Introduction
Epileptic seizures originate from diverse brain networks, exhibiting highly individualised, time-varying interictal and ictal patterns that often require uniquely personalised treatments to achieve complete seizure control. Around 1 in 100 people will be diagnosed with epilepsy (Vaughan et al., 2018), and approximately one-third of cases are resistant to medication (Kalilani et al., 2018). Many of these patients are candidates for surgery, which is a safe and effective treatment option to reduce or altogether eliminate seizures (Wiebe et al., 2001). This approach entails identifying and removing focal brain regions implicated in seizure generation, known as the epileptogenic zone. The seizure onset zone (SOZ) is the primary clinically identifiable proxy for these critical areas (Jehi, 2018). However, identifying the SOZ requires intracranial recordings of spontaneously occurring seizures. Seizures are comparatively rare events and often require extended hospital stays to capture, incurring high costs and stress for patients. Even with the current gold standard evaluation, outcomes remain comparatively poor in cases requiring invasive intracranial evaluation. Moreover, a putative SOZ is not identified at all in approximately 15% of cases, preventing surgical intervention (UK Children’s Epilepsy Surgery Collaboration et al., 2021).
@@
SPES elicits two primary types of responses: early and delayed. Early responses, emerging within 100ms post-stimulation and time-locked to the stimulus, are termed cortico-cortical evoked potentials (CCEPs) (Figure LABEL:fig:paradigmsD). These CCEPs signify directed or effective connectivity; for instance, if site A is stimulated, a CCEP observed at site B indicates directed neural propagation from A to B. While CCEPs occur in epileptogenic and non-epileptogenic tissue, their characteristics reflect the proximity of the stimulation and recording sites to the SOZ (Hays et al., 2023). Given their time-locked nature, CCEPs are typically averaged across 10-50 trials to enhance their signal-to-noise ratio.
In contrast, delayed responses, which occur between 100ms and 1s post-stimulation and resemble epileptiform discharges, reflect focal cortical irritability or even epileptogenicity (Figure LABEL:fig:paradigmsB). These responses have been reported in patients with temporal and frontal lobe epilepsies, and the removal of sites exhibiting these is associated with overall favourable outcomes (Valentín et al., 2005). Unlike the consistent and time-locked nature of CCEPs, delayed responses are neither consistent across trials nor time-locked to the stimulus, meaning they are less apparent when averaging across trials. Furthermore, other forms of inter-trial variability may also reflect epileptogenicity (Cornblath et al., 2023).
@@
Recent advances in machine learning, particularly in deep learning, are well-positioned to model complex multidimensional time-series data, such as those generated by SPES. This capability makes deep learning methods particularly promising for integrating multi-channel SPES responses to localise the SOZ. However, determining the most effective approach to leverage these advanced techniques remains a significant challenge due to particular data features, such as patient-specific electrode placements and variability in the background EEG activity.
@@
Secondly, effectively modelling patient-specific and heterogeneous multi-channel data presents a significant challenge. Most machine learning applications to time-series data assume a fixed number of channels with consistent ordering. However, intracranial EEG data varies in channel numbers and placements across patients, impeding model generalisability. Therefore, developing methods that can accommodate the unique and complex nature of patient-specific intracranial EEG datasets is an essential issue that needs to be addressed.
This paper introduces a CNN Transformer with cross-channel attention aimed at enhancing generalisability to unseen patients and managing the complexities of intracranial EEG data. Our study focuses on:
@@
  • Implementing and evaluating CNN Transformer models, outperforming existing methods with an AUROC of 0.730. We investigate the relationship between model performance and surgical outcome.
Generalizable Insights about Machine Learning in the Context of Healthcare
There is an increasing emphasis on identifying interictal biomarkers of the epileptogenic zone. One such approach is through stimulation techniques used to probe the epileptic brain, and deep learning is well suited to model the resulting multidimensional time-series data. In this work, we consider Single-Pulse Electrical Stimulation, focusing on two aspects: how these multi-channel responses to SPES can be represented and the best way to model heterogeneous intracranial electrode numbers and placements.
Our findings underscore the potential of deep learning to improve diagnostic approaches and treatment strategies in epilepsy. Specifically, we demonstrate the model’s capability to generalise to unseen patients within the same centre, setting a foundation for future research to extend this capability across diverse clinical environments and electrode configurations. This work holds broader implications for neurological research, particularly in contexts where precise localisation of brain activity is essential.
@@ 2.1 Localising the SOZ from SPES responses
Malone et al. (2022) evaluated various machine learning algorithms for localising the SOZ from SPES responses, classifying single-channel responses based on whether the recording electrode was in the SOZ. The most effective methods included a 1D convolutional neural network, a gradient-boosted decision tree, and an ensemble approach. In a different study, Yang et al. (2024) focused on computing various CCEP features to classify both stimulated and recording electrodes as SOZ/non-SOZ. These methods, however, consider each channel in isolation, leading to multiple predictions per channel and thereby not integrating the multidimensional data features into the classification. In contrast, Johnson et al. (2022) adopted a multi-channel approach using a divergent paradigm with a ResNet convolutional neural network. This method classifies the stimulated stereo-EEG (SEEG) channel based on responses from other channels. A limitation of this approach is that the network architecture chosen for this analysis requires a consistent number of channels and associates parameters with specific channel positions. To mitigate this, the authors randomly selected a subset of 40 channels for each pass, sorting these channels based on their proximity to the stimulated electrode. This method’s generalisability was validated using leave-one-out cross-validation. While this approach is suited to extracting temporal features, its effectiveness could benefit from adopting a more sophisticated method to focus on important response channels. Rather than having parameters associated with each of a fixed number of channels, capturing a context-aware representation across all channels may deal better with the heterogeneity of patient-specific seizure networks. Furthermore, this fixed channel count necessitates balancing the model’s prediction accuracy and generalisability.
@@ 3.1 Notation
To describe each network, we use the following notation. The output for a given stimulation trial is denoted as $x_{d}\in\mathbb{R}^{N\times C_{d,i}\times T}$ for the divergent paradigm, and $x_{c}\in\mathbb{R}^{N\times C_{c,i}\times T}$ for the convergent paradigm. Here, $N$ represents the number of trials, $C_{d,i}$ (or $C_{c,i}$ for a convergent paradigm) indicates the number of channels for patient $i$, and $T$ is the number of time points. Averaging these responses across trials results in $\bar{x}\in\mathbb{R}^{C_{i}\times T}$, which enhances the signal-to-noise ratio and facilitates a single prediction per electrode.
@@ 3.2 Models
    CNN Transformer
    An encoder-only Transformer (Vaswani et al., 2017). A CNN is applied to each channel, and cross-channel attention extracts features across channels. A convergent paradigm was used.
@@
Figure 2: Network architectures. Left: CNN. A ResNet creates a cross-channel embedding, and a multi-layer perceptron (MLP) classifies this as being in or out of the Seizure Onset Zone (SOZ). Right: CNN Transformer. First, a CNN is independently applied to each channel to extract per-channel features. These embeddings are then subjected to cross-channel attention before an MLP is used to classify the electrode as SOZ/non-SOZ.
@@ 3.2.3 CNN Transformer
Next, a learnable [CLS] token is prepended to these embeddings, and Transformer layers are applied with each channel treated as a token. An MLP is attached to the final [CLS] representation to classify the input as SOZ or non-SOZ. This approach enables feature extraction across an arbitrary number of channels, ensuring its suitability for intracranial EEG.
@@ 4.1 Dataset
For this study, we utilised an open-source dataset (van Blooijs et al., 2023a), which comprises electrocorticography (ECoG) recordings from 74 patients aged between 4 and 51 years undergoing evaluation for epilepsy surgery [^3] https://openneuro.org/datasets/ds004080. We restricted our analyses to the 35 patients for whom each electrode was labelled as within or outside the SOZ. The ECoG data were recorded at a sampling rate of 2048 Hz. Stimulation was performed at 0.2 Hz with a current intensity of 4mA or 8mA.
A total of 2066 ECoG electrodes were used, averaging 59 per patient. We restricted our analyses to stimulated electrodes to facilitate a fair comparison between both paradigms. Among these electrodes, 298 (14.4%) were identified as within the SOZ. The proportion of electrodes in the SOZ was relatively consistent across lobes (Figure LABEL:fig:lobe_distribution).
[…]
fig:lobe_distribution Refer to caption
@@ 4.2 Dataset split
We employed repeated k-fold cross-validation with $k=5$ and 5 repeats. To assess model generalisability to new patients and ECoG electrode locations, each patient was present in one fold, ensuring no patient data overlap between the training and testing sets. Further details are provided in Appendix A.
@@ 4.3 Preprocessing
Data were first bandpass filtered to 1-150 Hz in line with previous work in the same task (Johnson et al., 2022). Further preprocessing was applied in line with previous work using the same dataset (van Blooijs et al., 2023b), as outlined in Appendix B. These preprocessed time series were used as input to each model, as outlined in section 3.
@@ 4.4 Evaluation criteria
To assess model performance across repeated k-fold cross-validation, we employed both threshold-independent and threshold-dependent metrics, each computed per patient and then averaged across patients, folds, and repeats. We used the Area Under the Receiver Operating Characteristic (AUROC) as the threshold-independent metric. We report the specificity, sensitivity, and Youden’s index (Youden, 1950) for threshold-dependent metrics, which consider binary rather than continuous predictions. We test whether the CNN Transformer demonstrates an improvement in performance compared to CNNconvergent. The reported p-value assesses the statistical significance of the observed performance increase. Further details are included in Appendix C.
@@ 4.5 Hyperparameter tuning
We undertook a hyperparameter search for each model using Optuna (Akiba et al., 2019) to ensure that observed differences in model performance were not due to inadequate hyperparameter configurations. This process led to the selection of 49 channels for CNNdivergent and 37 channels for CNNconvergent. Further details of the hyperparameter search spaces and the chosen hyperparameters for each model are in Appendix D.
@@ 4.6 Channel sensitivity analysis
Although the rationale for employing cross-channel attention is its potential to extract features from relevant channels more effectively, it is worth considering the extent to which the CNN Transformer’s performance could be attributed to the availability of more channels during inference. If this were the case, it would indicate that the CNN Transformer benefits solely from handling arbitrary numbers of channels rather than the ability of cross-channel attention to extract relevant features. To investigate this aspect, we undertook a sensitivity analysis with the CNN Transformer, focusing on how its performance depends on the number of channels available during inference.
4.7 Model performance by outcome
An important consideration is whether the models offer complementary information to the SOZ. One possible mechanism for the models is that rather than approximating the SOZ – a proxy for the epileptogenic zone (EZ) – they may learn to approximate the EZ directly. Removal of the SOZ does not guarantee seizure freedom, so such a scenario could feasibly lead to improved outcomes. In general, we may expect that patients who became seizure-free had SOZs aligned closer with the EZ than those who continued to have seizures. Therefore, if the model approximates the EZ directly, we expect it to perform better in patients who became seizure-free. We tested this by performing an initial investigation to see the relationship between the CNN Transformer performance and surgical outcome. For this analysis, a modified version of the CNN Transformer that included the distance between stimulating and recording electrodes as a feature was used (mean AUROC: 0.745). We produced box plots of AUROC scores for seizure-free (Engel Class I) and non-seizure-free patients to investigate this. We then performed a one-tailed Mann-Whitney U test between these two groups to test for improved model performance in the seizure-free subcohort.
Table 1: Model Performance Metrics. Mean and standard deviation of the AUROC, AUPRC, specificity, sensitivity, and Youden’s index for each model across all test folds and runs.
Model AUROC Specificity Sensitivity Youden
CNNdivergent $0.574\pm 0.065$ $0.632\pm 0.118$ $0.478\pm 0.169$ $0.109\pm 0.106$
CNNconvergent $0.666\pm 0.067$ $0.692\pm 0.147$ $0.531\pm 0.199$ $0.223\pm 0.095$
CNN Transformer (ours) $0.730\pm 0.082$ $0.731\pm 0.130$ $0.589\pm 0.218$ $0.319\pm 0.134$